OpenStudy (anonymous):

Samantha wants to use her savings of $1150 to buy shirts and watches for her family. The total price of the shirts she bought was $84. The watches cost $99 each. What is the maximum number of watches that Samantha can buy with her savings @surjithayer

OpenStudy (anonymous):

@JBFlyBoiTv read my inbox I sent u and ill help also

OpenStudy (anonymous):

1150-84 --------- 99

OpenStudy (anonymous):

whole number = # of watches that can be purchased, decimal value is change left over in her pocket

OpenStudy (anonymous):

Let x be the no. of watches,then \[84+99x \le1150\] solve for x,the maximum integral value in this constraint is the required value.

OpenStudy (***[isuru]***):

price for the total no.of shirts + price for total no. of watches = the total amount of savings price for the total no.of shirts + price for total no. of watches = 1150 84 + price for total no. of watches = 1150 price for total no. of watches = 1150 - 84 = 1066 now divide 1066 by the price of a watch to get the no. of watches... 1066/99 = 10 .7676 so... the max. no .of watches Samantha can buy is 10